The volume of a cube with an edge length of u units is
u3 cubic units. If the length of
each edge is doubled, how many times larger is the volume of the new cube than the volume of the original cube?

Answer by ankor@dixie-net.com(12701)

The volume of a cube with an edge length of u units is u^3 cubic units.
If the length of each edge is doubled, how many times larger is the volume
of the new cube than the volume of the original cube?
:
Doubling the edge = 2u
Volume = (2u)^3 = 2^3 * u^3 = 8u^3
:
Volume is 8 times original cube